### v1

# Monotonicity of Recurrence in Random Walks

Abstract.
We consider non-homogeneous random walks on the positive quadrant in two dimensions. In the 1960’s the following question was asked: is it true if such a random walk $X$ is recurrent and $Y$ is another random walk that at every point is more likely to go down and more likely to go left than $Y$, then $Y$ is also recurrent?
We provide an example showing that the answer is negative. We also show that if either the random walk $X$ or $Y$ is sufficiently homogeneous then the answer is in fact positive.
1. Introduction
For the non-homogeneous reflecting random walk on $\mathbb{N}$ with transitions $p_{n}$ from $n$ to $n+1$ and $q_{n}=1-p_{n}$ from $n$ to $n-1$ (for $n>0$) there is an explicit characterization of when the random walk is recurrent/transient (see [2, Section XV.8]). From this it follows directly that if the random walk is recurrent and we consider a different one in which for all $n$ the $p_{n}^{\prime}$ satisfy $p_{n}^{\prime}\leq p_{n}$ then the new random walk is also recurrent, while if the first was transient while the second satisfies $p_{n}^{\prime}\geq p_{n}$ for all $n$ then the second is also transient. These monotonicity properties can also be easily seen via a coupling argument.
In the early 1960’s the question as to whether or not a similar monotonicity property holds for two dimensional random walks circulated among the probabilists at Princeton University. To the best of our knowledge this question has not been resolved to date. The purpose of the following note is to resolve this question.
We consider non-homogeneous random walks on the positive quadrant in two dimensions. Such a random walk is a Markov chain where the state space is $\mathbb{N}^{2}$ and one may only move to one of the four states immediately left, right, up, or down from the current state. We refer readers to [4] for a broad overview of random walks.
Our main interest is in understanding if recurrence is a monotone property with respect to the following partial order:
Definition 1.1.
For two random walks $X$ and $Y$ on $\mathbb{N}^{2}$, say $X\preceq Y$, or equivalently $Y\succeq X$, if for each state, the probability of going down is weakly greater in $X$ than in $Y$, the probability of going left is weakly greater in $X$ than in $Y$, the probability of going up is weakly greater in $Y$ than in $X$, and the probability of going right is weakly greater in $Y$ than in $X$.
Intuitively, $X\preceq Y$ means $X$ goes down and left more than $Y$.
The main question we address in the paper is the following:
Question 1.2.
Let $X$ and $Y$ be two non-homogeneous irreducible random walks on $\mathbb{N}^{2}$ with $X\preceq Y$. Is it true that if $Y$ is recurrent then so is $X$? In other words is it true that if $X$ is transient then so is $Y$?
Note that for two irreducible Markov chains $X$ and $Y$ on the same state space, if there only exist finitely many states for which the outgoing transition probabilities are different for $X$ and $Y$, i.e., $X$ and $Y$ agree on a cofinite set of states, then they will have the same recurrence behavior: $X$ is recurrent, positive recurrent, or transient if and only if $Y$ is recurrent, positive recurrent, or transient, respectively. Thus, when considering our question of monotonicity of recurrence, we can generalize Definition 1.1, as well as all other definitions that follow in this paper, so that the required inequalities only need to hold on a cofinite set of states. However, for sake of clarity and brevity, we henceforth do not discuss this equivalence up to a cofinite set.
The intuition behind this question is that if $Y$ is recurrent then it returns to the origin with probability $1$. As $X\preceq Y$ it is perhaps natural to expect that $X$ will also return to the origin with probability $1$, as indeed is the case in dimension one.
In our main result we prove that, in fact, the answer is negative.
Theorem 1.3.
There exists two non-homogeneous random walks $X\preceq Y$ such that $Y$ is positive recurrent and $X$ is transient. Moreover $X$ and $Y$ are elliptic, meaning all possible transitions have positive probabilities.
On the other hand, we do find conditions under which monotonicity preserves recurrence.
Definition 1.4.
Say a random walk on $\mathbb{N}^{2}$ is inward-homogeneous if it is of the following form:
$\begin{cases}\text{right}&\text{w.p. }r_{o}\\ \text{up}&\text{w.p. }u_{o},\end{cases}\qquad\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{x}\\ \text{right}&\text{w.p. }r_{x}\\ \text{up}&\text{w.p. }u_{x}\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }d_{y}\\ \text{up}&\text{w.p. }u_{y}\\ \text{right}&\text{w.p. }r_{y},\end{cases}\qquad\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{q}\\ \text{right}&\text{w.p. }r_{q}\\ \text{up}&\text{w.p. }u_{q}\\ \text{down}&\text{w.p. }d_{q},\end{cases}$
for $(0,0)$, $\{(i,0):i\geq 1\}$, $\{(0,j):j\geq 1\}$, and $\{(i,j):i,j\geq 1\}$, respectively, where
$r_{o}\geq r_{x}=r_{q},\quad\ell_{x}\geq\ell_{q},\quad u_{x}\geq u_{q},\quad u_{o}\geq u_{y}=u_{q},\quad d_{y}\geq d_{q},\quad\text{and}\quad r_{y}\geq r_{q}.$
One can view an inward-homogeneous random walk as being given by the transition probabilities $\ell_{q}$, $r_{q}$, $u_{q}$, and $d_{q}$ from the general quadrant case, along with three stochastic “redirection rules” if a forbidden move is selected, i.e., moving down while on the positive $x$-axis, moving left while on the positive $y$-axis, or moving left or down while at the origin. The only condition for inward-homogeneity is that the positive $x$-axis cannot redirect towards the right, and the positive $y$-axis cannot redirect up.
We now show that under an inward-homogeneity condition, recurrence and positive recurrence are monotonic properties with respect to $\preceq$.
Theorem 1.5.
If an irreducible inward-homogeneous random walk $X$ on $\mathbb{N}^{2}$ is recurrent (respectively, positive recurrent), then any irreducible random walk $Y$ on $\mathbb{N}^{2}$ such that $Y\preceq X$ is also recurrent (respectively, positive recurrent).
For the reverse direction, where we want transience of $X$ to imply transience of $Y$ for $Y\succeq X$, we can of course take the contrapositive of Theorem 1.5, but this requires $Y$ to be inward-homogeneous. The following theorem instead assumes $X$ is inward-homogeneous, and its proof is very similar to that of Theorem 1.5.
Theorem 1.6.
If an irreducible inward-homogeneous random walk $X$ on $\mathbb{N}^{2}$ is transient, then any irreducible random walk $Y$ on $\mathbb{N}^{2}$ such that $Y\succeq X$ is also transient.
Using Theorems 1.5 and 1.6 and their contrapositives, we obtain the following result that allows for application of monotonicity when neither random walk is inward-homogeneous, though an inward-homogeneous random walk must exist between the two random walks.
Corollary 1.7.
Let $X$, $Y$, and $Z$ be irreducible random walks on $\mathbb{N}^{2}$ where $Y$ is inward-homogeneous and $X\preceq Y\preceq Z$. If $Z$ is recurrent then $X$ is recurrent, and if $X$ is transient then $Z$ is transient.
A similar monotonicity question can be asked for slabs, i.e., sets of the form $\{(i,j)\in\mathbb{Z}^{2}:i\geq 0,0\leq j\leq k\}$, which can be viewed as an interpolation between the one-dimensional environment of non-homogeneous random walks on $\mathbb{N}$, for which monotonicity always holds trivially, and the two-dimensional environment that we study. We obtain analogous results for slabs in Section 4.
Rayleigh’s monotonicity law, introduced by Rayleigh [3] in 1899, may be able to make some statements on this question of when monotonicity of recurrence occurs, but is severely limited by the requirement that both $X$ and $Y$ must be Markov chains that correspond to electric networks (see [1] for a comprehensive overview of Rayleigh’s monotonicity law). A Markov chain corresponds to an electric network if there exist constants $c_{xy}\geq 0$ for every unordered pair of two distinct states $x$ and $y$, such that letting $c_{x}=\sum_{y}c_{xy}$ for all states $x$, assuming only finitely many terms in each sum are positive, we have the transition probability from $x$ to $y$ is $c_{xy}/c_{x}$ and the transition probability from $y$ to $x$ is $c_{xy}/c_{y}$. When the Markov chain is positive recurrent, this is equivalent to the Markov chain being reversible, where the stationary distribution $\pi$ is given by $\pi_{x}\propto c_{x}$. This imposes certain conditions on a random walk on $\mathbb{N}^{2}$: for example, for 4 states $w$, $x$, $y$, and $z$, we have $\frac{c_{wx}}{c_{xy}}\cdot\frac{c_{xy}}{c_{yz}}\cdot\frac{c_{yz}}{c_{wz}}\cdot\frac{c_{wz}}{c_{wx}}=1$. Letting $(w,x,y,z)=((0,0),(1,0),(1,1),(0,1))$ in an inward-homogeneous random walk on $\mathbb{N}^{2}$, this implies $\frac{\ell_{x}d_{q}r_{y}u_{o}}{u_{x}\ell_{q}d_{y}r_{o}}=1$. Considering two other “squares” $((1,0),(2,0),(2,1),(1,1))$ and $((0,1),(1,1),(1,2),(0,2))$ yields two additional conditions $\frac{r_{x}\ell_{q}}{\ell_{x}r_{q}}=1$ and $\frac{u_{q}d_{y}}{u_{y}d_{q}}=1$. These conditions severely limit the applicability of Rayleigh monotonicity to two random walks $X\preceq Y$.
In Section 2, we prove Theorem 1.3. In Section 3, we prove Theorems 1.5 and 1.6. Lastly, in Section 4, we prove analogues of Theorems 1.3, 1.5, 1.6 and 1.7 for slabs.
2. A Counterexample to Monotonicity
In this section, we show that recurrence, positive recurrence, and transience are not monotonic properties of random walks on $\mathbb{N}^{2}$ with respect to the partial order $\preceq$. To do so, we construct a positive recurrent random walk $X$ and a transient random walk $Y$ such that $Y\preceq X$, proving Theorem 1.3.
See Figure 1 for a visual schematic of the constructions for $X$ and $Y$. For each random walk, the transition probabilities for the states on the positive $x$-axis are the same, as well as for the states on the positive $y$-axis, and for the states not on either axis; in that sense, there are only four types of vertices, and we draw the expectation of their transition for each of the four types.
(a) The positive recurrent random walk $X$.
(b) The transient random walk $Y$.
Figure 1. Schematic diagrams of the constructed positive recurrent $X$ and transient $Y$ used to prove Theorem 1.3. Arrows show the expected movement starting from the origin, $x$-axis, $y$-axis, and the rest of the quadrant.
As one can see, the behavior on the $X$-axis is transient while the behavior on the $Y$-axis is recurrent. For the recurrent walk in the interior of the quadrant, there is a drift to the left while for the transient walk there is a drift downwards. This is the idea behind the following example.
2.1. Recurrent Random Walk
Consider the following random walk on $\mathbb{N}^{2}$:
$\begin{cases}\text{right}&\text{w.p. }\frac{9}{14}\\ \text{up}&\text{w.p. }\frac{5}{14},\end{cases}\qquad\qquad\begin{cases}\text{up}&\text{w.p. }\frac{1}{2}\\ \text{right}&\text{w.p. }\frac{5}{12}\\ \text{left}&\text{w.p. }\frac{1}{12}\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }\frac{89}{120}\\ \text{up}&\text{w.p. }\frac{5}{24}\\ \text{right}&\text{w.p. }\frac{1}{20},\end{cases}\qquad\qquad\begin{cases}\text{up}&\text{w.p. }\frac{3}{8}\\ \text{down}&\text{w.p. }\frac{3}{8}\\ \text{left}&\text{w.p. }\frac{5}{24}\\ \text{right}&\text{w.p. }\frac{1}{24},\end{cases}$
for $(0,0)$, $\{(i,0):i\geq 1\}$, $\{(0,j):j\geq 1\}$, and $\{(i,j):i,j\geq 1\}$, respectively. It is easy to see that this Markov chain is irreducible and 2-periodic. Hence, to show that it is positive recurrent, it suffices to prove a stationary distribution exists. It is straightforward to verify that the following distribution is stationary:
$\pi_{i,j}=\begin{cases}\frac{1}{Z}\cdot\left(\frac{5}{7}\right)^{i+j}&i,j\geq 1\\ \frac{1}{Z}\cdot\frac{3}{4}\left(\frac{5}{7}\right)^{i}&j=0\text{ and }i\geq 1\\ \frac{1}{Z}\cdot\frac{5}{6}\left(\frac{5}{7}\right)^{j}&i=0\text{ and }j\geq 1\\ \frac{1}{Z}\cdot\frac{35}{72}&(i,j)=(0,0),\end{cases}$
where $Z<\infty$ normalizes $\pi$ so that $\sum_{i,j}\pi_{i,j}=1$.
2.2. Transient Random Walk
Now, consider the following random walk on $\mathbb{N}^{2}$:
$\begin{cases}\text{right}&\text{w.p. }\frac{9}{14}\\ \text{up}&\text{w.p. }\frac{5}{14},\end{cases}\qquad\qquad\begin{cases}\text{up}&\text{w.p. }\frac{49}{100}\\ \text{right}&\text{w.p. }\frac{41}{100}\\ \text{left}&\text{w.p. }\frac{1}{10}\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }\frac{19}{25}\\ \text{up}&\text{w.p. }\frac{5}{25}\\ \text{right}&\text{w.p. }\frac{1}{25},\end{cases}\qquad\qquad\begin{cases}\text{up}&\text{w.p. }\frac{1}{100}\\ \text{down}&\text{w.p. }\frac{74}{100}\\ \text{left}&\text{w.p. }\frac{21}{100}\\ \text{right}&\text{w.p. }\frac{4}{100},\end{cases}$
for $(0,0)$, $\{(i,0):i\geq 1\}$, $\{(0,j):j\geq 1\}$, and $\{(i,j):i,j\geq 1\}$, respectively. Again, this Markov chain is irreducible and 2-periodic. Note that it strictly satisfies the desired partial order relationship with the Markov chain from Section 2.1 in the sense that for $\{(i,j):i,j\geq 1\}$, the left and down probabilities strictly increased while the right and up probabilities strictly decreased, and similarly for the axes, except of course the down probability and left probability for the $x$-axis and $y$-axis, respectively, must be zero and thus do not increase.
To show this Markov chain is transient, consider a random walk $X_{n}$ with $X_{0}=(0,0)$. It suffices to show
$\sum_{k=1}^{\infty}\mathbb{P}(X_{k}=(0,0))=\sum_{k=1}^{\infty}\mathbb{P}(X_{2k}=(0,0))<\infty,$
as this is the expected number of times $X_{n}$ returns to $(0,0)$.
Let $S_{k}=x_{2k}-y_{2k}$ where $X_{n}=(x_{n},y_{n})$. Let $A_{k}=S_{k}-S_{k-1}$ for $k\geq 1$. Regardless of $X_{2k-2}$, which is in $\{(i,j):i+j\text{ is even}\}$, we can couple $A_{k}$ with $B_{k}$ such that $A_{k}\geq B_{k}$, where $B_{k}$ has marginal distribution
$B_{k}=\begin{cases}-2&\text{w.p. }\frac{1668}{10000}\\ 0&\text{w.p. }\frac{6651}{10000}\\ 2&\text{w.p. }\frac{1681}{10000},\end{cases}$
for all $k\geq 1$. As $A_{k}\in\{-2,0,2\}$ always, this simply means that regardless of $X_{2k-2}$, we have $A_{k}=-2$ with probability at most $\frac{1668}{10000}$ and have $A_{k}=2$ with probability at least $\frac{1681}{10000}$; this can be verified for all possible values of $X_{2k-2}$. This coupling is tight, i.e., $A_{k}=B_{k}$, when $y_{2k-2}=0$ and $x_{2k-2}>0$. Crucially, the $B_{k}$ are mutually independent. Let $T_{k}=\sum_{\ell=1}^{k}B_{\ell}\leq\sum_{\ell=1}^{k}A_{\ell}=S_{k}$. Note that $T_{k}$ has a positive drift and by Hoeffding’s inequality, for all $k\geq 1$,
$\mathbb{P}(X_{2k}=(0,0))$ $\leq\mathbb{P}(S_{k}=0)\leq\mathbb{P}(T_{k}\leq 0)\leq\exp\left(-\frac{2\left(\frac{13}{10000}k\right)^{2}}{16k}\right)=\exp\left(\frac{-169k}{80000}\right),$
so
$\sum_{k=1}^{\infty}\mathbb{P}(X_{2k}=(0,0))\leq\sum_{k=1}^{\infty}\exp\left(\frac{-169k}{80000}\right)<\infty.$
3. Monotonicity in Homogeneous Random Walks
We now consider conditions under which recurrence and transience are indeed monotonic properties with respect to the partial order $\preceq$ and prove Theorem 1.5.
We will use the following notation. Let $\leq$ denote the partial order on $\mathbb{Z}^{2}$ given by $(i_{1},j_{1})\leq(i_{2},j_{2})$ if $i_{1}\leq i_{2}$ and $j_{1}\leq j_{2}$. For two elements $x=(i_{1},j_{1})$ and $y=(i_{2},j_{2})$ in $\mathbb{Z}^{2}$, we let $x+y$ denote $(i_{1}+i_{2},j_{1}+j_{2})$. We view $\mathbb{N}^{2}$ as a subset of $\mathbb{Z}^{2}$
Proof of Theorem 1.5.
Let $X$ and $Y$ both start at the origin. Note that irreducibility of $X$ implies it is elliptic, i.e., all probabilities in Definition 1.4 are positive. It suffices to show that we can couple $X$ and $Y$ such that $Y_{t}\leq X_{t}+s_{t}$ for some $s_{t}\in\{(2,0),(1,1),(0,2)\}$ for all $t$. Let $S=\{(2,0),(1,1),(0,2)\}$. This is because for each time $t$ such that $X_{t}=(0,0)$, we have $Y_{t}\in\{(0,0),(2,0),(1,1),(0,2)\}$, as $X$ and $Y$ are both 2-periodic, so the probability that $Y_{t+2}=(0,0)$ is at least some positive constant $p>0$. So if $X$ is recurrent and thus $X_{t}=(0,0)$ for infinitely many $t$ regardless of $X_{0}$, the probability that $Y$ never returns to $(0,0)$ is 0. And if $X$ is positive recurrent so that some upper bound $\tau<\infty$ is greater than or equal to the four expected hitting times of $(0,0)$ starting at one of the four states in $S\cup\{(0,0)\}$, then the expected return time for $Y$ is at most $\frac{\tau+2}{p}$, and thus $Y$ is also positive recurrent.
As $Y\preceq X$, we can view each step of $Y$ as following a two-stage process: at the current state, select a direction based on the transition probabilities of $X$ at this state, and then if the direction is up or right, with some probability we instead go left or down. The second “inward shift” step converts the transition probabilities of $X$ at this state to the desired transition probabilities of $Y$ at this state.
Using the redirection framework for inward-homogeneity, we can subdivide each step of $X$ with another two-stage process: first selecting left, right, up, or down based on $\ell_{q}$, $r_{q}$, $u_{q}$, and $d_{q}$, and then redirecting as necessary. This subdivision transforms the process of determining each step of $Y$ into a three-stage process.
We can now couple $X$ and $Y$ as follows, so that the first two stages of $Y$ coincide with the two stages of $X$. While $X$ and $Y$ may not be at the same state, we first use the same selection of left, right, up, or down based on $\ell_{q}$, $r_{q}$, $u_{q}$, and $d_{q}$. Then $X$ and $Y$ both redirect as necessary, but if $X$ and $Y$ use the same redirection rule, i.e., both are on the positive $x$-axis, both are on the positive $y$-axis, or both are at the origin, the same redirection choice is made. If they use different redirection rules, the choices can be made arbitrarily. Finally, if the selected move for $Y$ is up or right, we inward shift as necessary.
It now remains to verify that this coupling satisfies $Y_{t}\leq X_{t}+s_{t}$ for some $s_{t}\in S$ for all $t$, which we prove by induction. We start with $X_{0}=Y_{0}=(0,0)$, and for fixed $t\geq 1$, suppose that $Y_{t-1}\leq X_{t-1}+s_{t-1}$ for some $s_{t-1}\in S$. Let $Y_{t}^{\prime}$ be the state obtained from $Y_{t}$ by ignoring inward shifting at timestep $t$; it suffices to show $Y_{t}^{\prime}\leq X_{t}+s_{t}$ for some $s_{t}\in S$, as $Y_{t}\leq Y_{t}^{\prime}$ always. The only way $Y_{t}^{\prime}\leq X_{t}+s_{t-1}$ could not hold is if $X$ and $Y$ have different redirection rules (with no redirection being a fourth possibility), as otherwise $Y_{t}^{\prime}-Y_{t-1}=X_{t}-X_{t-1}$ and thus $Y_{t}^{\prime}\leq X_{t}+s_{t-1}$ then follows from the inductive hypothesis.
We split into cases depending on the selection of left, right, up, or down based on $\ell_{q}$, $r_{q}$, $u_{q}$, and $d_{q}$, i.e., the first stage of the three-stage coupling. Note that up and right movements never get redirected, so we only need to check down and left. Let $X_{t}=(i_{t}^{(1)},j_{t}^{(1)})$ and let $Y_{t}=(i_{t}^{(2)},j_{t}^{(2)})$. Also let $Y_{t}^{\prime}=(i_{t}^{(2^{\prime})},j_{t}^{(2^{\prime})})$. Note that the 2-periodicity of our random walks implies
(3.1) $i_{t}^{(1)}+j_{t}^{(1)}\equiv i_{t}^{(2)}+j_{t}^{(2)}\equiv i_{t}^{(2^{\prime})}+j_{t}^{(2^{\prime})}\equiv t\pmod{2}.$
Case 1: left. Redirections only occur on the $y$-axis. If both $X$ and $Y^{\prime}$ redirect but under different rules, then one of them is at the origin while the other is on the positive $y$-axis. If $Y_{t-1}=(0,0)$, then $Y_{t}^{\prime}\leq(1,1)\leq X_{t}+(1,1)$. Otherwise, $X_{t-1}=(0,0)$ and $Y_{t-1}=(0,2)$, where $Y^{\prime}$ can only move down or right; if $Y^{\prime}$ moves right, then $Y_{t}^{\prime}-X_{t}=(1,1)$ or $Y_{t}^{\prime}-X_{t}=(0,2)$ if $X$ moves up or right, respectively, and if $Y^{\prime}$ moves down, then $Y_{t}^{\prime}\leq X_{t}+(1,1)$.
If instead exactly one of $X$ and $Y^{\prime}$ redirects, first suppose $X$ redirects while $Y^{\prime}$ moves left, so that $i_{t-1}^{(1)}=0$ and $i_{t-1}^{(2)}\in\{1,2\}$. If $i_{t-1}^{(2)}=2$ so that $i_{t}^{(2^{\prime})}=1$, we have $s_{t-1}=(2,0)$ so $j_{t-1}^{(2)}\leq j_{t-1}^{(1)}$, and thus $Y_{t}^{\prime}\leq X_{t}+(1,1)$. Otherwise $i_{t-1}^{(2)}=1$ so that $i_{t}^{(2^{\prime})}=0$, and thus $s_{t-1}\in\{(1,1),(2,0)\}$ so $j_{t-1}^{(2)}\leq j_{t-1}^{(1)}+1$, and thus $Y_{t}^{\prime}\leq X_{t}+(0,2)$.
Lastly, suppose $Y^{\prime}$ redirects while $X$ moves left, so that $i_{t-1}^{(2)}=0$ and $i_{t-1}^{(1)}\geq 1$. If $Y^{\prime}$ moves up, meaning $Y_{t-1}=(0,0)$, then $Y_{t}^{\prime}\leq X_{t}+(0,2)$. As we always have $j_{t-1}^{(2)}\leq j_{t-1}^{(1)}+2$, if $Y^{\prime}$ moves down, then $Y_{t}^{\prime}\leq X_{t}+(0,2)$. If $Y^{\prime}$ moves right, then $Y_{t}^{\prime}\leq X_{t}+(0,2)$ if $i_{t-1}^{(1)}\geq 2$, and if $i_{t-1}^{(1)}=1$, by (3.1) we have $j_{t-1}^{(2)}\leq j_{t-1}^{(1)}+1$, so $Y_{t}^{\prime}\leq X_{t}+(1,1)$.
Case 2: down. The arguments for this case are the same as for case 1, except reflected across the line $y=x$. ∎
Remark 3.1.
The assumptions of irreducibility in Theorem 1.5 can be removed, rephrasing recurrence as recurrence of the state $(0,0)$. The proof of Theorem 1.5 only uses irreducibility to imply ellipticity, which is used to show that if $Y_{t}\in\{(0,0),(2,0),(1,1),(0,2)\}$, then in each of these four cases we have $Y_{t+2}=(0,0)$ with positive probability. However, this follows from the assumptions that $Y\preceq X$ and $X$ is inward-homogeneous. If $Y_{t}=(0,0)$ implies $Y_{t+2}\neq(0,0)$ with probability 1, without loss of generality suppose $Y_{t+1}=(1,0)$ with positive probability. Then $Y$ cannot move left from $(1,0)$, implying $\ell_{q}=\ell_{x}=0$ for $X$, so with some probability $X_{1}=(1,0)$, from which it cannot return to $(0,0)$, contradicting recurrence of $X$. Essentially identical arguments address the cases $Y_{t}=(2,0)$ and $Y_{t}=(0,2)$. For $Y_{t}=(1,1)$, there are four cases for how one could have $Y_{t+2}\neq(0,0)$ with probability 1.
Case 1: $Y$ cannot move down from $(0,1)$ and cannot move left from $(1,0)$. Then $X$ also cannot make these moves, and thus cannot ever return to $(0,0)$, contradicting recurrence of $(0,0)$.
Case 2: $Y$ cannot move down from $(0,1)$ and cannot move down from $(1,1)$. Then $X$ cannot ever move down, yet as $Y_{t}=(1,1)$, with positive probability $X$ can reach $(1,1)$, from which it cannot return to $(0,0)$, contradicting recurrence of $(0,0)$.
Case 3: $Y$ cannot move left from $(1,1)$ and cannot move left from $(1,0)$. This case is the reflection of case 2, and essentially the same argument holds.
Case 4: $Y$ cannot move left from $(1,1)$ and cannot move down from $(1,1)$. Then $X$ cannot move left or down in $\{(i,j):i,j\geq 1\}$, yet as $Y_{t}=(1,1)$, with positive probability $X$ can reach $\{(i,j):i,j\geq 1\}$, from which it cannot return to $(0,0)$, contradicting recurrence of $(0,0)$.
We now prove Theorem 1.6.
Proof of Theorem 1.6.
Let $X$ and $Y$ both start at the origin. It suffices to show that we can couple $X$ and $Y$ such that $Y_{t}\geq X_{t}-s_{t}$ for some $s_{t}\in\{(2,0),(1,1),(0,2)\}$ for all $t$. Let $S=\{(2,0),(1,1),(0,2)\}$. This is because $X$ being transient implies that with some positive probability, $X$ visits $\{(i,j):i+j\leq 2\}$ finitely often, and thus $Y$ returns to $(0,0)$ finitely often.
We use the same coupling as in the proof of Theorem 1.5, except in the third stage, if the selected move for $Y$ is down or left, we may apply an “outward shift” and instead go up or right. We prove $Y_{t}\geq X_{t}-s_{t}$ for some $s_{t}\in S$ for all $t$ by induction on $t$. We start with $X_{0}=Y_{0}=(0,0)$, and for fixed $t\geq 1$, suppose $Y_{t-1}\geq X_{t-1}-s_{t-1}$ for some $s_{t-1}\in S$. Let $Y_{t}^{\prime}$ be the state obtained from $Y_{t}$ by ignoring outward shifting at timestep $t$; it suffices to show $Y_{t}^{\prime}\geq X_{t}-s_{t}$ for some $s_{t}\in S$, as $Y_{t}\geq Y_{t}^{\prime}$ always. Similar to before, we only need to consider when $X$ and $Y$ have different redirection rules, and we split into cases depending on the selection of left, right, up, or down. As up and right movements never get redirected, by reflecting across the line $y=x$, without loss of generality it suffices to address the left case. Let $X_{t}=(i_{t}^{(1)},j_{t}^{(1)})$, $Y_{t}=(i_{t}^{(2)},j_{t}^{(2)})$, and $Y_{t}^{\prime}=(i_{t}^{(2^{\prime})},j_{t}^{(2^{\prime})})$.
As redirections only occur on the $y$-axis, if both $X$ and $Y^{\prime}$ redirect but under different rules, then one of them is at the orign while the other is on the positive $y$-axis. If $X_{t-1}=(0,0)$, then $Y_{t}^{\prime}\geq(1,1)-(1,1)\geq X_{t}-(1,1)$. Otherwise $Y_{t-1}=(0,0)$, so $X_{t-1}=(0,2)$, where $X$ can only move down or right; if $X$ moves right, then $X_{t}-Y_{t}^{\prime}=(1,1)$ or $X_{t}-Y_{t}^{\prime}=(0,2)$ if $Y^{\prime}$ moves up or right, respectively, and if $X$ moves down, then $Y_{t}^{\prime}\geq X_{t}-(1,1)$.
If instead exactly one of $X$ and $Y^{\prime}$ redirects, first suppose $Y$ redirects while $X^{\prime}$ moves left, so that $i_{t-1}^{(2)}=0$ and $i_{t-1}^{(1)}\in\{1,2\}$. If $i_{t-1}^{(1)}=2$ so that $i_{t}^{(1)}=1$, we have $s_{t-1}=(2,0)$ so $j_{t-1}^{(2)}\geq j_{t-1}^{(1)}$, and thus $Y_{t}^{\prime}\geq X_{t}-(1,1)$. Otherwise $i_{t-1}^{(1)}=1$ so that $i_{t}^{(1)}=0$, and thus $s_{t-1}\in\{(1,1),(2,0)\}$ so $j_{t-1}^{(2)}\geq j_{t-1}^{(1)}-1$, and thus $Y_{t}^{\prime}\geq X_{t}-(0,2)$.
Lastly, suppose $X$ redirects while $Y^{\prime}$ moves left, so that $i_{t-1}^{(1)}=0$ and $i_{t-1}^{(2)}\geq 1$. If $X$ moves up, meaning $X_{t-1}=(0,0)$, then $Y_{t}^{\prime}\geq(0,-1)=X_{t}-(0,2)$. As we always have $j_{t-1}^{(2)}\geq j_{t-1}^{(1)}-2$, if $X$ moves down, then $Y_{t}^{\prime}\geq X_{t}-(0,2)$. If $X$ moves right, then $Y_{t}^{\prime}\geq X_{t}-(0,2)$ if $i_{t-1}^{(2)}\geq 2$, and if $i_{t-1}^{(2)}=1$, by (3.1) we have $j_{t-1}^{(2)}\geq j_{t-1}^{(1)}-1$, so $Y_{t}^{\prime}\geq X_{t}-(1,1)$. ∎
4. Random Walks on a Slab
For positive integer $k$, let the slab of thickness $k$ be $\{(i,j)\in\mathbb{Z}^{2}:i\geq 0,0\leq j\leq k\}$. Intuitively, the slab is an intermediate environment for random walks between the 1-dimensional random walks on $\mathbb{N}$ and the 2-dimensional random walks on $\mathbb{N}^{2}$, which are recovered in the cases $k=0$ and $k=\infty$, respectively. For a slab of thickness $k$, we partition the slab into six (possibly empty) regions: the center $\{(i,j):i\geq 1,0<j<k\}$, the upper boundary $\{(i,k):i\geq 1\}$, the lower boundary $\{(i,0):i\geq 1\}$, the left boundary $\{(0,j):0<j<k\}$, the origin $(0,0)$, and the upper corner $(0,k)$.
In Section 4.1, similar to Section 2, we provide a counterexample to monotonicity of recurrence, positive recurrence, and transience of random walks on the slab with respect to the partial order $\preceq$ defined analogously to Definition 1.1. The construction works for all $k\geq 2$, and while a different construction can work for $k=1$, for sake of brevity we do not study the case $k=1$. Furthermore, in Section 4.2, similar to Section 3, we show that for $k\geq 2$, monotonicity holds under a different and arguably more natural partial order of random walks on the slab, assuming a homogeneity condition similar to Definition 1.4 of inward-homogeneity. Recalling the definition of inward-homogeneity is based on transition probabilities $\ell_{q}$, $r_{q}$, $u_{q}$, and $d_{q}$ in the general case of states lying on neither axis, this concept of homogeneity does not neatly transfer to the case $k=1$, where there is no center. Thus, we henceforth restrict attention to $k\geq 2$.
4.1. A counterexample to monotonicity
Using the essentially identical definition of $\preceq$ for random walks on $\mathbb{N}^{2}$ from Definition 1.1 for random walks on the slab, we first show that recurrence, positive recurrence, and transience are not monotonic properties of random walks on the slab with respect to the partial order $\preceq$. In this sense, random walks on the slab are similar to random walks on $\mathbb{N}^{2}$, and our construction of a positive recurrent random walk $X$ and a transient random walk $Y$ such that $Y\preceq X$ follows the same intuition. For our random walk on $\mathbb{N}^{2}$, the positive $x$-axis moves outwards while the positive $y$-axis moves inwards and the primary change from $X$ to $Y$ is changing the center from predominantly moving towards the $y$-axis to predominantly moving towards the $x$-axis. For our slab construction, we preserve the role of the positive $x$-axis, i.e., the lower boundary, and use the upper boundary in place of the positive $y$-axis.
However, our proof technique for the recurrent side is different. Fix positive integer $k\geq 2$, and consider the following random walk on the slab of thickness $k$:
$\begin{cases}\text{up}&\text{w.p. }0.65\\ \text{right}&\text{w.p. }0.33\\ \text{left}&\text{w.p. }0.01\\ \text{down}&\text{w.p. }0.01,\end{cases}\qquad\qquad\begin{cases}\text{right}&\text{w.p. }0.52\\ \text{up}&\text{w.p. }0.47\\ \text{left}&\text{w.p. }0.01,\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }0.49\\ \text{left}&\text{w.p. }0.48\\ \text{right}&\text{w.p. }0.03,\end{cases}$
$\begin{cases}\text{up}&\text{w.p. }0.97\\ \text{right}&\text{w.p. }0.02\\ \text{down}&\text{w.p. }0.01,\end{cases}\qquad\qquad\begin{cases}\text{right}&\text{w.p. }0.99\\ \text{up}&\text{w.p. }0.01,\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }0.50\\ \text{right}&\text{w.p. }0.50,\end{cases}$
for the center, lower boundary, upper boundary, left boundary, origin, and upper corner, respectively. It is easy to see that this Markov chain is irreducible and 2-periodic. To be explicit about the 2-periodicity, we partition the slab of thickness $k$ into the two sets $T_{0}=\{(i,j)\in\mathbb{Z}^{2}:i\geq 0,0\leq j\leq k,i+j\text{ is even}\}$ and $T_{1}=\{(i,j)\in\mathbb{Z}^{2}:i\geq 0,0\leq j\leq k,i+j\text{ is odd}\}$, where starting in a state in $T_{0}$ we must move to a state in $T_{1}$, and vice versa. The precise argument to show positive recurrence depends on the parity of $k$. We first assume $k$ is even, where the argument is slightly simpler.
We claim that it suffices to show that for any starting state in $T_{0}$ within $k$ distance of at least one of the two corners, i.e., the origin and upper corner, the expected hitting time of $\{(0,0),(0,k)\}$, meaning the expected time to reach either of the two corners, not including time 0 if starting at one of the two corners, is finite. If this holds, then let $r<\infty$ be greater than or equal to all of these expected hitting times. There exists some probability $p>0$ such that with at least probability $p$, we reach the upper corner after exactly $k$ steps, starting at either of the two corners. It suffices to show that the expected return time of the upper corner is finite, but we can bound our expected return time by $\frac{r+k}{p}<\infty$, by considering each time we return to one of the two corners, which takes at most $r$ time in expectation, and then whether or not we reach the upper corner after $k$ more steps, which occurs with probability $p$; if not, we are at one of the starting states in the definition of $r$, and we try again, where each trial takes at most $r+k$ time in expectation and has success probability at least $p$.
Now, consider starting at a state in $T_{0}$ within $k$ distance of at least one of the two corners, and let our random walk be $X_{t}$. Let $\tau$ be the hitting time of $\{(0,0),(0,k)\}$, so that we must show $\mathbb{E}[\tau]<\infty$. However, modify $X_{t}$ such that if we return to either of the two corners, we stop, i.e., for all $t\geq\tau$, we have $X_{t}=X_{\tau}$. Similar to the proof of transience in Section 2.2, let $S_{n}=x_{2n}-y_{2n}$ where $X_{t}=(x_{t},y_{t})$. As $X_{2n}\in T_{0}$ for all nonnegative integers $n$, it is straightforward, though slightly tedious, to verify that $\mathbb{E}[S_{n+1}-S_{n}|S_{n}]\leq-0.0892$ regardless of the value of $S_{n}\in T_{0}$, for $n\geq 1$. This inequality is tight when $S_{n}$ is on the lower boundary. Crucially, without the modification of $X_{t}$ such that we stop upon returning to either of the two corners, this inequality would be violated if $S_{n}\in\{(0,0),(0,k)\}$. We have $X_{0}=(x_{0},y_{0})$ is fixed, with $S_{n}\geq-k$ for all $n$, so
$-k-x_{0}+y_{0}\leq\lim_{n\to\infty}\mathbb{E}[S_{n}-S_{0}].$
As $|S_{1}-S_{0}|\leq 2$ always, we have
$-k-x_{0}+y_{0}-2$ $\leq\lim_{n\to\infty}\mathbb{E}[S_{n}-S_{1}]=\sum_{n=1}^{\infty}\mathbb{E}[S_{n+1}-S_{n}]\leq\sum_{n=1}^{\infty}-0.0892\cdot\Pr(X_{2n}\not\in\{(0,0),(0,k)\})$
$=-0.0892\sum_{n=1}^{\infty}\Pr(\tau>2n)=0.0892-0.0446\cdot\mathbb{E}[\tau],$
where the last equality holds because $\tau$ must be a positive even integer. This bounds
$\mathbb{E}[\tau]\leq\frac{k+x_{0}-y_{0}+2}{0.0446}+2<\infty.$
The argument when $k$ is odd is nearly identical, with some adjustments due to the fact that $(0,k)$ lies in $T_{1}$ instead of $T_{0}$. Let $U=\{(0,k),(1,0),(0,1)\}$. Applying essentially the same argument as when $k$ is even, it suffices to show that for any starting state in $T_{1}$ within $k+1$ distance of at least one of the points in $U$, the expected hitting time of $U$ is finite. If we let $r<\infty$ be an upper bound on these expected hitting times and let $p$ be a lower bound on the probability of reaching $(0,k)$ after exactly $k+1$ steps, starting at one of the three points in $U$, then we can bound our expected return time of the upper corner by $\frac{r+k+1}{p}<\infty$.
Now, consider starting at a state in $T_{1}$ within $k+1$ distance of at least one of the points in $U$, and let this random walk be $X_{t}$, modified so that if we return to $U$, we stop. Let $\tau$ be the hitting of $U$. Again letting $S_{n}=x_{2n}-y_{2n}$, we continue to have $\mathbb{E}[S_{n+1}-S_{n}|S_{n}]\leq-0.0892$ regardless of the value of $S_{n}\in T_{1}$, for $n\geq 1$. Again, the modification to stop upon reaching $U$ is critical for this inequality to work, as the inequality would be violated if we moved from either of the corners, which cannot occur for $n\geq 1$ by the stopping rule, though it can occur for $n=0$ if we started in $U$. The same argument as before yields $\mathbb{E}[\tau]\leq\frac{k+x_{0}-y_{0}+2}{0.0446}+2<\infty$, which completes the proof of positive recurrence.
Our proof technique for the transient random walk on the slab is identical to that on $\mathbb{N}^{2}$ in Section 2.2. Consider the following random walk on the slab of thickness $k$:
$\begin{cases}\text{up}&\text{w.p. }0.01\\ \text{right}&\text{w.p. }0.32\\ \text{left}&\text{w.p. }0.02\\ \text{down}&\text{w.p. }0.65,\end{cases}\qquad\qquad\begin{cases}\text{right}&\text{w.p. }0.51\\ \text{up}&\text{w.p. }0.46\\ \text{left}&\text{w.p. }0.03,\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }0.50\\ \text{left}&\text{w.p. }0.49\\ \text{right}&\text{w.p. }0.01,\end{cases}$
$\begin{cases}\text{up}&\text{w.p. }0.01\\ \text{right}&\text{w.p. }0.01\\ \text{down}&\text{w.p. }0.98,\end{cases}\qquad\qquad\begin{cases}\text{right}&\text{w.p. }0.99\\ \text{up}&\text{w.p. }0.01,\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }0.51\\ \text{right}&\text{w.p. }0.49,\end{cases}$
for the center, lower boundary, upper boundary, left boundary, origin, and upper corner, respectively. Again, this Markov chain is irreducible and 2-periodic, and strictly satisfies the desired partial order relationship with the positive recurrent Markov chain in the sense that the left and down probabilities strictly increased while the right and up probabilities strictly decreased, whenever they are allowed to be positive. We use the same notation of $S_{n}$, $A_{n}$, and $B_{n}$ from Section 2.2, where $B_{n}$ has marginal distribution
$B_{n}=\begin{cases}-2&\text{w.p. }0.2401\\ 0&\text{w.p. }0.4998\\ 2&\text{w.p. }0.2601.\end{cases}$
The $-2$ probability is tight when on the upper boundary, and the $2$ probability is tight when on the lower boundary. As $\mathbb{E}[B_{k}]>0$, our Hoeffding argument from Section 2.2 implies this random walk is transient.
4.2. Monotonicity in Homogeneous Random Walks
As the slab has finite thickness, there is no reason to believe that increasing the down probabilities will increase the chance that the random walk is recurrent; by reflecting across $y=k/2$, the up and down probabilities play identical roles. And as seen in Section 4.1, increasing the down probabilities can allow for counterexamples of monotonicity. To this end, we introduce a more natural definition for a partial order on random walks on the slab.
Definition 4.1.
For two random walks $X$ and $Y$ on the slab of thickness $k$, say $X\trianglelefteq Y$, or equivalently $Y\trianglerighteq X$, if for each state, the probability of going left is weakly greater in $X$ than in $Y$, the probability of going right is weakly greater in $Y$ than in $X$, the probability of going up is the same in $X$ as in $Y$, and the probability of going down is the same in $X$ as in $Y$.
Intuitively, $X\trianglelefteq Y$ means $X$ goes left more than $Y$. Similar to Definition 1.4, we introduce a definition of a homogeneous random walk on the slab such that monotonicity holds.
Definition 4.2.
Say a random walk on the slab is homogeneous if it is of the following form:
$\begin{cases}\text{left}&\text{w.p. }\ell_{q}\\ \text{right}&\text{w.p. }r_{q}\\ \text{up}&\text{w.p. }u_{q}\\ \text{down}&\text{w.p. }d_{q},\end{cases}\qquad\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{x}\\ \text{right}&\text{w.p. }r_{x}\\ \text{up}&\text{w.p. }u_{x},\end{cases}\qquad\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{u}\\ \text{right}&\text{w.p. }r_{u}\\ \text{down}&\text{w.p. }d_{u},\end{cases}$
$\begin{cases}\text{down}&\text{w.p. }d_{y}\\ \text{up}&\text{w.p. }u_{y}\\ \text{right}&\text{w.p. }r_{y},\end{cases}\qquad\qquad\begin{cases}\text{right}&\text{w.p. }r_{o}\\ \text{up}&\text{w.p. }u_{o},\end{cases}\qquad\qquad\begin{cases}\text{right}&\text{w.p. }r_{c}\\ \text{down}&\text{w.p. }d_{c},\end{cases}$
for the center, lower boundary, upper boundary, left boundary, origin, and upper corner, respectively, where
$r_{c}\geq r_{u}\geq r_{q}\leq r_{x}\leq r_{o},\quad r_{y}\geq r_{q},\quad\ell_{u}\geq\ell_{q}\leq\ell_{x},\quad d_{u},d_{y},d_{c}\geq d_{q},\quad\text{and}\quad u_{x},u_{y},u_{o}\geq u_{q}.$
Similar to an inward-homogeneous random walk on $\mathbb{N}^{2}$, we can view a homogeneous random walk on the slab as being given by the four transition probabilities from the general center case, along with redirection rules if a forbidden move is selected. Within this redirection framework, the only nontrivial inequalities are $r_{c}\geq r_{u}$ and $r_{o}\geq r_{x}$, which can be interpreted as the existence of a coupling of the upper corner and boundary redirection rules such that if both attempt to move up and the upper boundary redirects to the right, then the upper corner must redirect to the right as well, and symmetrically for the origin and lower boundary.
[…]
Theorem 4.3.
For positive integer $k\geq 2$, if an irreducible homogeneous random walk $X$ on the slab of thickness $k$ is recurrent (respectively, positive recurrent), then any irreducible random walk $Y$ on the slab of thickness $k$ such that $Y\trianglelefteq X$ is also recurrent (respectively, positive recurrent).
Proof.
Our proof is structurally similar to that of Theorem 1.5, so for sake of brevity and clarity we will omit additional elaboration when the reasoning is essentially identical to before. Let $X$ and $Y$ both start at the origin. Let $X_{t}=(i_{t}^{(1)},j_{t}^{(1)})$ and $Y_{t}=(i_{t}^{(2)},j_{t}^{(2)})$. It suffices to show that we can couple $X$ and $Y$ such that $|j_{t}^{(2)}-j_{t}^{(1)}|+\max\{i_{t}^{(2)}-i_{t}^{(1)},0\}\leq 2\lceil k/2\rceil$ for all $t$. We couple $X$ and $Y$ in the same manner as before, where inward shifting only ever changes a right movement in $Y$ to a left movement. We couple the upper corner and boundary redirection rules as described in Definition 4.2, which implies that if one of $X$ and $Y$ is on the upper boundary and the other is at the upper corner, then ignoring inward shifting, if they (both) attempt to move up and the one on the upper boundary redirects right, the one at the upper corner will redirect right as well. We similarly couple the origin and lower boundary redirection rules.
We use the same notation of $Y_{t}^{\prime}=(i_{t}^{(2^{\prime})},j_{t}^{(2^{\prime})})$ as before. It suffices to show that for fixed $t\geq 1$, if $|j_{t-1}^{(2)}-j_{t-1}^{(1)}|+\max\{i_{t-1}^{(2)}-i_{t-1}^{(1)},0\}\leq 2\lceil k/2\rceil$, then $|j_{t}^{(2^{\prime})}-j_{t}^{(1)}|+\max\{i_{t}^{(2^{\prime})}-i_{t}^{(1)},0\}\leq 2\lceil k/2\rceil$, as inward shifting weakly decreases the second term. If neither $X$ nor $Y$ redirect, the desired inequality trivially holds by induction. Now, suppose there is exactly one redirection; we divide into casework on who redirects and the direction that the random walks attempted to go in.
Case 1: $X$ redirects while attempting to go up. Then $Y$ goes up, decreasing the first term by 1 from time $t-1$ to $t$, so the inequality holds regardless of how $X$ redirects, as its movement can only change one term by at most 1.
Case 2: $X$ redirects while attempting to go down. This case is symmetrically equivalent to case 1.
Case 3: $X$ redirects while attempting to go left. Then $Y$ goes left, decreasing the second term by 1, so the inequality holds regardless of how $X$ redirects.
Case 4: $Y$ redirects while attempting to go up. This case follows from the same argument as in case 1.
Case 5: $Y$ redirects while attempting to go down. This case is symmetrically equivalent to case 4.
Case 6: $Y$ redirects while attempting to go left. This implies $i_{t}^{(2^{\prime})}-i_{t}^{(1)}\leq 0$, so we trivially bound $|j_{t}^{(2^{\prime})}-j_{t}^{(1)}|+\max\{i_{t}^{(2^{\prime})}-i_{t}^{(1)},0\}\leq k+0\leq 2\lceil k/2\rceil$.
Finally, suppose there are two different redirections, meaning $X$ and $Y$ are in two different non-central regions; we divide into casework on the direction that the random walks attempted to go in. Note that we can ignore any cases where the redirection rules pick the same direction, as then the two terms are unchanged.
Case 1: $X$ and $Y$ redirect while attempting to go up. If $X$ is to the right, by 2-periodicity we bound the first term by 1 and the second term by 0. If $Y$ is to the right, meaning $X$ is at the upper corner, the situation is more complicated. If $Y$ redirects down, then (ignoring $X$ moving down as then their movements will be the same) $X$ moves right, so the first term increases by 1 while the second term decreases by 1, so the inequality holds by induction. If $Y$ redirects left then the second term decreases by 1, so the inequality holds regardless of how $X$ redirects. If $Y$ redirects right, our coupling implies $X$ redirects right, which implies the inequality continues to hold.
Case 2: $X$ and $Y$ redirect while attempting to go down. This case is symmetrically equivalent to case 1.
Case 3: $X$ and $Y$ redirect while attempting to go left. As $X$ and $Y$ are both on the $y$-axis, by 2-periodicity they are an even distance apart. We have $\max\{i_{t-1}^{(2)}-i_{t-1}^{(1)},0\}=0$. If $k$ is odd, or if $k$ is even and they are at most $k-2$ apart, then $|j_{t-1}^{(2)}-j_{t-1}^{(1)}|\leq 2\lceil k/2\rceil-2$ and the sum of the two terms can increase by at most 2, yielding the desired inequality. Otherwise, $X$ and $Y$ are on the two corners. If either moves vertically then the first term decreases by 1 so the inequality holds regardless of how the other redirects; otherwise, they both move right, in which case the inequality continues to hold. ∎
[…]
Theorem 4.4.
For positive integer $k\geq 2$, if an irreducible homogeneous random walk $X$ on the slab of thickness $k$ is transient, then any irreducible random walk $Y$ on the slab of thickness $k$ such that $Y\trianglerighteq X$ is also transient.
Proof.
Following the reasoning of the proof of Theorem 1.6 and using the notation of the proof of Theorem 4.3, it suffices to couple $X$ and $Y$ such that $|j_{t}^{(1)}-j_{t}^{(2)}|+\max\{i_{t}^{(1)}-i_{t}^{(2)},0\}\leq 2\lceil k/2\rceil$ for all $t$. The coupling used is the same as that in the proof of Theorem 4.3, where inward shifting is replaced by outward shifting as in the proof of Theorem 1.6. The inductive verification that the coupling satisfies the desired inequality is identical to that in the proof of Theorem 4.3, where in all the casework we swap the roles of $X$ and $Y$. ∎
Analogous to Corollary 1.7, the following corollary of Theorems 4.3 and 4.4 allows for application of monotonicity when neither random walk is homogeneous, though a homogeneous random walk must exist between the two random walks.
Corollary 4.5.
For positive integer $k\geq 2$, let $X$, $Y$, and $Z$ be irreducible random walks on the slab of thickness $k$ where $Y$ is homogeneous and $X\trianglelefteq Y\trianglelefteq Z$. If $Z$ is recurrent then $X$ is recurrent, and if $X$ is transient then $Z$ is transient.
Acknowledgements
We thank Jason Liu for helpful discussions.